\begin{thebibliography}{10}

\bibitem{barkhagen2018stochastic}
{\sc Mathias Barkhagen, Ngoc~Huy Chau, {E}ric Moulines, Mikl{o}s R{a}sonyi,
  Sotirios Sabanis, and Ying Zhang}, {\em On stochastic gradient langevin
  dynamics with dependent data streams in the logconcave case}, arXiv preprint
  arXiv:1812.02709,  (2018).

\bibitem{bierkens2019zig}
{\sc Joris Bierkens, Paul Fearnhead, Gareth Roberts, et~al.}, {\em The zig-zag
  process and super-efficient sampling for bayesian analysis of big data}, The
  Annals of Statistics, 47 (2019), pp.~1288--1320.

\bibitem{bou2020coupling}
{\sc Nawaf Bou-Rabee, Andreas Eberle, Raphael Zimmer, et~al.}, {\em Coupling
  and convergence for hamiltonian monte carlo}, Annals of Applied Probability,
  30 (2020), pp.~1209--1250.

\bibitem{bouchard2018bouncy}
{\sc Alexandre Bouchard-Cote, Sebastian~J Vollmer, and Arnaud Doucet}, {\em The
  bouncy particle sampler: A nonreversible rejection-free markov chain monte
  carlo method}, Journal of the American Statistical Association, 113 (2018),
  pp.~855--867.

\bibitem{brooks2011handbook}
{\sc Steve Brooks, Andrew Gelman, Galin Jones, and Xiao-Li Meng}, {\em Handbook
  of markov chain monte carlo}, CRC press, 2011.

\bibitem{chau2019stochastic}
{\sc Ngoc~Huy Chau, Eric Moulines, Miklos Rasonyi, Sotirios Sabanis, and Ying
  Zhang}, {\em On stochastic gradient langevin dynamics with dependent data
  streams: the fully non-convex case}, arXiv preprint arXiv:1905.13142,
  (2019).

\bibitem{chen2021stochastic}
{\sc Yongxin Chen, Tryphon~T Georgiou, and Michele Pavon}, {\em Stochastic
  control liaisons: Richard sinkhorn meets gaspard monge on a schrodinger
  bridge}, SIAM Review, 63 (2021), pp.~249--313.

\bibitem{cheng2018convergence}
{\sc Xiang Cheng and Peter Bartlett}, {\em Convergence of langevin mcmc in
  kl-divergence}, Proceedings of Machine Learning Research, Volume 83:
  Algorithmic Learning Theory,  (2018), pp.~186--211.

\bibitem{cheng2018sharp}
{\sc Xiang Cheng, Niladri~S Chatterji, Yasin Abbasi-Yadkori, Peter~L Bartlett,
  and Michael~I Jordan}, {\em Sharp convergence rates for langevin dynamics in
  the nonconvex setting}, arXiv preprint arXiv:1805.01648,  (2018).

\bibitem{dai1991stochastic}
{\sc Paolo Dai~Pra}, {\em A stochastic control approach to reciprocal diffusion
  processes}, Applied mathematics and Optimization, 23 (1991), pp.~313--329.

\bibitem{dalalyan2017further}
{\sc Arnak~S Dalalyan}, {\em Further and stronger analogy between sampling and
  optimization: Langevin monte carlo and gradient descent}, arXiv: Statistics
  Theory,  (2017).

\bibitem{dalalyan2017theoretical}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em Theoretical
  guarantees for approximate sampling from smooth and log-concave densities},
  Journal of the Royal Statistical Society: Series B (Statistical Methodology),
  79 (2017), pp.~651--676.

\bibitem{dalalyan2019user-friendly}
{\sc Arnak~S Dalalyan and Avetik~G Karagulyan}, {\em User-friendly guarantees
  for the langevin monte carlo with inaccurate gradient}, Stochastic Processes
  and their Applications, 129 (2019), pp.~5278--5311.

\bibitem{dunson2020hastings}
{\sc David~B Dunson and JE~Johndrow}, {\em The hastings algorithm at fifty},
  Biometrika, 107 (2020), pp.~1--23.

\bibitem{durmus2016high-dimensional}
{\sc Alain Durmus and Eric Moulines}, {\em High-dimensional bayesian inference
  via the unadjusted langevin algorithm.}, arXiv: Statistics Theory,  (2016).

\bibitem{durmus2016sampling}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em Sampling from a
  strongly log-concave distribution with the unadjusted langevin algorithm},
  arXiv: Statistics Theory,  (2016).

\bibitem{durmus2017nonasymptotic}
{\sc Alain Durmus, Eric Moulines, et~al.}, {\em Nonasymptotic convergence
  analysis for the unadjusted langevin algorithm}, The Annals of Applied
  Probability, 27 (2017), pp.~1551--1587.

\bibitem{eldan2020}
{\sc Ronen Eldan, Joseph Lehec, Yair Shenfeld, et~al.}, {\em Stability of the
  logarithmic sobolev inequality via the follmer process}, in Annales de
  l'Institut Henri Poincar{e}, Probabilit{e}s et Statistiques, vol.~56,
  Institut Henri Poincar{e}, 2020, pp.~2253--2269.

\bibitem{follmer1985}
{\sc Hans Follmer}, {\em An entropy approach to the time reversal of diffusion
  processes}, in Stochastic Differential Systems Filtering and Control,
  Springer, 1985, pp.~156--163.

\bibitem{follmer1986}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em Time reversal on
  wiener space}, in Stochastic processes-mathematics and physics, Springer,
  1986, pp.~119--129.

\bibitem{follmer1988}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em Random fields and
  diffusion processes}, in Ecole dEte de Probabilites de Saint-Flour XV--XVII,
  1985--87, Springer, 1988, pp.~101--203.

\bibitem{gelfand1990sampling}
{\sc Alan~E Gelfand and Adrian~FM Smith}, {\em Sampling-based approaches to
  calculating marginal densities}, Journal of the American statistical
  association, 85 (1990), pp.~398--409.

\bibitem{geman1984stochastic}
{\sc Stuart Geman and Donald Geman}, {\em Stochastic relaxation, gibbs
  distributions, and the bayesian restoration of images}, IEEE Transactions on
  pattern analysis and machine intelligence,  (1984), pp.~721--741.

\bibitem{hale2010asymptotic}
{\sc Jack~K Hale}, {\em Asymptotic behavior of dissipative systems}, no.~25,
  American Mathematical Soc., 2010.

\bibitem{hastings1970monte}
{\sc W~Keith Hastings}, {\em Monte carlo sampling methods using markov chains
  and their applications},  (1970).

\bibitem{sfs21}
{\sc Jian Huang, Yuling Jiao, Lican Kang, Xu~Liao, Jin Liu, and Yanyan Liu},
  {\em Schr{o}dinger-follmer sampler: Sampling without ergodicity}, arXiv
  preprint arXiv: 2106.10880,  (2021).

\bibitem{lehec2013representation}
{\sc Joseph Lehec}, {\em Representation formula for the entropy and functional
  inequalities}, in Annales de l'IHP Probabilit{e}s et statistiques, vol.~49,
  2013, pp.~885--899.

\bibitem{leonard2014survey}
{\sc Christian L{e}onard}, {\em A survey of the schr{o}dinger problem and some
  of its connections with optimal transport}, Discrete Continuous Dynamical
  Systems-A, 34 (2014), p.~1533.

\bibitem{liu2008monte}
{\sc Jun~S Liu}, {\em Monte Carlo strategies in scientific computing}, Springer
  Science Business Media, 2008.

\bibitem{ma2019sampling}
{\sc Yi-An Ma, Yuansi Chen, Chi Jin, Nicolas Flammarion, and Michael~I Jordan},
  {\em Sampling can be faster than optimization}, Proceedings of the National
  Academy of Sciences, 116 (2019), pp.~20881--20885.

\bibitem{menz2014poincare}
{\sc Georg Menz, Andr{e} Schlichting, et~al.}, {\em Poincar{e} and logarithmic
  sobolev inequalities by decomposition of the energy landscape}, Annals of
  Probability, 42 (2014), pp.~1809--1884.

\bibitem{metropolis1953equation}
{\sc Nicholas Metropolis, Arianna~W Rosenbluth, Marshall~N Rosenbluth,
  Augusta~H Teller, and Edward Teller}, {\em Equation of state calculations by
  fast computing machines}, The journal of chemical physics, 21 (1953),
  pp.~1087--1092.

\bibitem{mou2019improved}
{\sc Wenlong Mou, Nicolas Flammarion, Martin~J Wainwright, and Peter~L
  Bartlett}, {\em Improved bounds for discretization of langevin diffusions:
  Near-optimal rates without convexity}, arXiv preprint arXiv:1907.11331,
  (2019).

\bibitem{raginsky2017non}
{\sc Maxim Raginsky, Alexander Rakhlin, and Matus Telgarsky}, {\em Non-convex
  learning via stochastic gradient langevin dynamics: a nonasymptotic
  analysis}, in Conference on Learning Theory, PMLR, 2017, pp.~1674--1703.

\bibitem{robert2010introducing}
{\sc Christian~P Robert, George Casella, and George Casella}, {\em Introducing
  monte carlo methods with r}, vol.~18, Springer, 2010.

\bibitem{schrodinger1932theorie}
{\sc Erwin Schr{o}dinger}, {\em Sur la th{e}orie relativiste de l{e}lectron et
  l'interpr{e}tation de la m{e}canique quantique}, in Annales de l'institut
  Henri Poincar{e}, vol.~2, 1932, pp.~269--310.

\bibitem{Tierney1994Markov}
{\sc Luke Tierney}, {\em Markov chains for exploring posterior distributions},
  The Annals of Statistics, 22 (1994), pp.~1701--1728.

\bibitem{tzen2019theoretical}
{\sc Belinda Tzen and Maxim Raginsky}, {\em Theoretical guarantees for sampling
  and inference in generative models with latent diffusions}, in Conference on
  Learning Theory, PMLR, 2019, pp.~3084--3114.

\bibitem{wang2009log}
{\sc Feng-Yu Wang et~al.}, {\em Log-sobolev inequalities: different roles of
  ric and hess}, The Annals of Probability, 37 (2009), pp.~1587--1604.

\bibitem{zhang2019nonasymptotic}
{\sc Ying Zhang, Omer~Deniz Akyildiz, Theo Damoulas, and Sotirios Sabanis},
  {\em Nonasymptotic estimates for stochastic gradient langevin dynamics under
  local conditions in nonconvex optimization}, arXiv preprint arXiv:1910.02008,
   (2019).

\end{thebibliography}

\documentclass[final,a4paper]{siamltex}
\usepackage{amsmath,amssymb,amsfonts}
\usepackage{verbatim}
\usepackage{changepage}
\usepackage{bm,color,url}
\usepackage{algorithm,algorithmic}
\usepackage{geometry}
\usepackage{setspace}
\usepackage[notref,notcite]{showkeys}
\usepackage{hyperref}

\newtheorem{Lemma*}{theorem}
\newtheorem{remark}{Remark}[section]
\newtheorem{remark*}{Remark}

\newcommand{\bZ}{{\bf Z}}
\newcommand{\bI}{\mbox{\bf I}}

\def\Ebb{\mathbb{E}}

\title{Convergence Analysis of Schr{\"o}dinger-F{\"o}llmer Sampler without Convexity}
\author{Yuling Jiao
\thanks{School of Mathematics and Statistics, Wuhan University, Wuhan, China.
Email: yulingjiaomath@whu.edu.cn} \quad \and Lican Kang
\thanks{School of Mathematics and Statistics, Wuhan University, Wuhan, China.
Email: kanglican@whu.edu.cn}
\quad\and
Yanyan Liu
\thanks{School of Mathematics and Statistics, Wuhan University, Wuhan, China.
Email: liuyy@whu.edu.cn}
\quad\and Youzhou Zhou \thanks{ Department of Pure mathematics, Xi'an Jiaotong-Liverpool University, Suzhou, China.
Email: youzhou.zhou@xjtlu.edu.cn}
}

\begin{document}

\maketitle

\begin{abstract}
Schr\"{o}dinger-F\"{o}llmer sampler (SFS) \cite{sfs21},  is a novel  and efficient approach for sampling from possibly unnormalized distributions without ergodicity.  SFS is based on the Euler-Maruyama discretization of Schr\"{o}dinger-F\"{o}llmer diffusion process
$$\mathrm{d} X_{t}=-\nabla U\left(X_t, t\right) \mathrm{d} t+\mathrm{d} B_{t}, \quad t \in[0,1],\quad X_0=0$$
 on the unit interval, which transports the degenerate distribution at time zero to the target distribution at time one.
In \cite{sfs21}, the consistency of SFS is established under a restricted assumption that  
{the potential $U(x,t)$
is uniformly  (on $t$) strongly 
convex
(on $x$).
} In this paper we provide   a nonasymptotic error bound of SFS in  Wasserstein distance under  some smooth  and bounded conditions on the
 density ratio of the target distribution over the standard normal distribution,  but   without requiring   the strongly
{convexity of the
 potential.
}
\end{abstract}

\section{Introduction}\label{introduction}
Sampling from  possibly unnormalized distributions  is an  import task in Bayesian statistics and machine learning.
 Ever since the Metropolis-Hastings (MH) algorithm \cite{metropolis1953equation,hastings1970monte} was introduced, various random sampling methods were proposed, including
Gibbs sampler,  random walk sampler, independent sampler, Lagevin sampler,   bouncy particle sampler, zig-zag sampler  \cite{geman1984stochastic,gelfand1990sampling,Tierney1994Markov,liu2008monte,
robert2010introducing,bouchard2018bouncy,bierkens2019zig}, among others, see \cite{brooks2011handbook,dunson2020hastings}  and the references therein. The above mentioned sampling algorithms generate random samples   by running an ergodic Markov chain whose stationary distribution is the target distribution.

In \cite{sfs21}, the  Schr\"{o}dinger-F\"{o}llmer sampler (SFS),  a novel sampling approach  without requiring the property of ergodicity is proposed. SFS is  based on  the  Schr\"{o}dinger-F\"{o}llmer diffusion process, defined as
\begin{align}\label{sde}
\mathrm{d} X_{t}=b\left(X_{t}, t\right) \mathrm{d} t+\mathrm{d} B_{t}, \quad t \in[0,1],\quad X_0=0,
\end{align}
where the drift function $$b(x,t) = -\nabla U(x,t)=\frac{\Ebb_{Z\sim N\left(0,\bI_{p}\right)}[\nabla f(x+\sqrt{1-t}Z)]}{\Ebb_{Z\sim N\left(0,\bI_{p}\right)}[f(x+\sqrt{1-t} Z)]}:\mathbb{R}^p \times [0,1]\rightarrow \mathbb{R}^1$$
 with $f(\cdot)=\frac{d \mu }{d N\left(0,\bI_{p}\right)}(\cdot)$.
According to \cite{leonard2014survey} and \cite{eldan2020}, the process  $\{X_t\}_{t\in [0,1]}$  in (\ref{sde}) was first formulated by F\"{o}llmer \cite{follmer1985, follmer1986, follmer1988} when studying the Schr\"{o}dinger bridge problem \cite{schrodinger1932theorie}.
The main feature of the above  Schr\"{o}dinger-F\"{o}llmer  process is that  it interpolates $\delta_{0}$ and the target $\mu$ in time $[0,1]$, i.e.,  $X_1 \sim \mu$, see Proposition \ref{SBP}.  SFS  samples from $\mu$ via the  following Euler-Maruyama discretization of (\ref{sde}),
$$
Y_{t_{k+1}}=Y_{t_k}+ sb\left(Y_{t_k}, t_k\right)+\sqrt{s}\epsilon_{k+1},
~
Y_{t_0} = 0,~ k=0,1,\ldots, K-1,
$$
where $s = 1/K$ is the  step size, $t_k = s k$,  and $\{\epsilon_{k}\}_{k=1}^{K}$ are independent and identically distributed from $N(0,\bI_{p})$.
 If the expectations in the   drift term $b(x,t)$   do not have  analytical forms,  one can use   Monte Carlo  method   to evaluate
  $b\left(Y_{t_k}, t_k\right)$ approximately, i.e., one can sample from $\mu$ according
 $$
\widetilde{Y}_{t_{k+1}}=\widetilde{Y}_{t_k}+ s\tilde{b}_m\left(\widetilde{Y}_{t_k}, t_k\right)+\sqrt{s}\epsilon_{k+1},
~
\widetilde{Y}_{t_0} = 0, ~k=0,1,\ldots, K-1,$$
  where $\tilde{b}_m(\widetilde{Y}_{t_{k}},t_{k})=
\frac{\frac{1}{m}\sum_{j=1}^m[\nabla f(\widetilde{Y}_{t_{k}}+\sqrt{1-t_{k}}Z_j)]}{\frac{1}{m}\sum_{j=1}^{m} [f(\widetilde{Y}_{t_{k}}+\sqrt{1-t_{k}} Z_j)]}$ with $Z_1,...Z_m$ i.i.d $N(0,\bI_{p})$. The numerical simulations in   \cite{sfs21} demonstrate  that SFS outperforms the exiting samplers based on
egordicity.

In Section 4.2 of \cite{sfs21}, they prove that
$$W_2(\mbox{Law}(\widetilde{Y}_{t_K}),\mu)\rightarrow 0, \ \ \mathrm{as} \ \  s \rightarrow 0, m \rightarrow \infty$$ under a restricted assumption that
{the potential $U(x,t)$ is uniformly strongly convex, i.e.,
\begin{equation}\label{cond4}
U(x,t)-U(y,t)-\nabla U(y,t)^{\top}(x-y)\geq (M/2)\left\|x-y\right\|^2_2, \forall x,y\in \mathbb{R}^p,  \forall t\in [0,1],
\end{equation}
where $M$ is one finite and positive constant.
}
In this paper we provide a new analysis of the above SFS iteration. We establish a nonasymptotic error bound on
$W_2(\mbox{Law}(\widetilde{Y}_{t_K}),\mu)$ under   the condition that  $f$ and $\nabla f$ are Lipschitz continuous and $f$ has positive  lower bound,     but without using  the uniformly strongly convexity requirement (\ref{cond4}).

The rest of this paper is organized as follows. In Section \ref{method},  we  recall the SFS  method.
In Section \ref{Theorey}, we present our  theoretical analysis.
We conclude in Section \ref{conlusion}. Proofs for all the theorems are provided in Appendix \ref{append}.

\section{Schr{\"o}dinger-F{\"o}llmer sampler}\label{method}
In this section we  recall the Schr{\"o}dinger-F{\"o}llmer  sampler briefly. More   backgrounds  on the  Schr{\"o}dinger-F{\"o}llmer diffusion process please see \cite{dai1991stochastic,leonard2014survey,chen2021stochastic,sfs21}.

Let $\mu \in \mathcal{P}\left(\mathbb{R}^{p}\right)$ be the target distribution and absolutely continuous with respect to the $p$-dimensional standard normal measure $G=N(0,\bI_{p})$. Let $$f(x)=\frac{d\mu}{dG}(x).$$
We assume that
\begin{itemize}
\item[$(\textbf{A1})$] $f,\nabla f$ are Lipschitz continuous with constant $\gamma$,
\item[$(\textbf{A2})$] There exists $\xi>0$ such that $f\geq \xi$.
\end{itemize}
Define the heat semigroup $Q_t, t\in [0,1]$ as
$$
Q_{t} f(x)=\Ebb_{Z \sim G}[f(x+\sqrt{t} Z)].
$$

\begin{proposition}\label{SBP}
Define a drift function
$$
b(x, t)=\nabla \log Q_{1-t} f(x).
$$
If $f$ satisfies assumptions $(\textbf{A1})$ and $(\textbf{A2})$, then the Schr{\"o}dinger-F\"{o}llmer diffusion
\begin{align}\label{sch-equation}
\mathrm{d} X_{t}=b\left(X_{t}, t\right) \mathrm{d} t+\mathrm{d} B_{t}, \quad t \in[0,1], \quad X_{0}=0,
\end{align}
has a unique strong solution and $X_{1} \sim \mu$.
\end{proposition}
\begin{remark}
\begin{itemize}
{\item[(i)]
From the definition of $b(x,t)$ in Proposition \ref{SBP}, it follows that
$U(x,t)=-\log Q_{1-t}f(x)$.
}
\item[(ii)]
If the target distribution is $\mu(dx)=\exp (-V(x)) d x / C$ with the normalized constant $C$,
then $f(x)=\frac{\left(\sqrt{2\pi}\right)^p}{C}\exp (-V(x)+\frac{\|x\|_2^2}{2})$.
Once $V(x)$ is twice differentiable and
$$
\lim_{R\to\infty}\sup_{\|x\|_2\geq R}\exp \left(-V(x)+\|x\|_2^{2} / 2\right)\|x-\nabla V\|_2<\infty,
$$
$$
\lim_{R\to\infty}\sup_{\|x\|_2\geq R}\exp \left(-V(x)+\|x\|_2^{2} / 2\right)\|Hess(V)\|_2<\infty,
$$
then both $f$ and $\nabla f$ are Lipschitz continuous, i.e., (\textbf{A1}) holds.
(\textbf{A2}) is equivalent to the growth condition on the potential that  $V(x) \leq \frac{\|x\|^2}{2} - \log \xi + \mathrm{contant}$.
\item[(iii)] Under (\textbf{A1}) and (\textbf{A2}), some calculation shows that  $$\|\nabla f\|_2\leq \gamma, \|Hess(f)\|_2\leq\gamma,$$ and
$$
\sup_{x\in\mathbb{R}^p,t\in[0,1]}\|\nabla Q_{1-t}f(x)\|_2\leq \gamma,
\sup_{x\in\mathbb{R}^p,t\in[0,1]}\|Hess(Q_{1-t}f(x))\|_2\leq \gamma,
$$ and
$$
b(x,t)=\frac{\nabla Q_{1-t}f(x)}{Q_{1-t}f(x)}, ~\nabla b(x,t)=\frac{Hess(Q_{1-t}f)(x)}{Q_{1-t}f(x)}-b(x,t)b(x,t)^{\top}.
$$
We conclude that
$$
\sup_{x\in\mathbb{R}^p,t\in[0,1]}\|b(x,t)\|_2\leq \frac{\gamma}{\xi},
\sup_{x\in\mathbb{R}^p,t\in[0,1]}\|\nabla b(x,t)\|_2\leq \frac{\gamma}{\xi}+\frac{\gamma^2}{\xi^2}.
$$
\end{itemize}
\end{remark}
Proposition   \ref{SBP} shows  that the Schr{\"o}dinger-F\"{o}llmer diffusion will transport $\delta_0$ to the target $\mu $ on the unite time interval. 
Since  drift term $b(x,t)$ is scale-invariant with respect to $f$ in the sense that
$b(x, t)=\nabla \log Q_{1-t} Cf(x), \forall C>0$. Therefore, the Schr{\"o}dinger-F\"{o}llmer diffusion can be used for sampling from $\mu(dx)=\exp (-V(x)) d x/C$,  where  the normalizing constant of  $C$ may not to be known.
To this end, we use the  Euler-Maruyama method to discretize  the  Schr{\"o}dinger-F\"{o}llmer diffusion (\ref{sch-equation}).
Let $$t_{k}=k\cdot s, \ \ ~k=0,1,  \ldots, K, \ \ ~\mbox{with}~ \ \ s=1 / K, \ \ Y_{t_{0}}=0,$$
the Euler-Maruyama scheme reads
\begin{align}\label{emd}
Y_{t_{k+1}}=Y_{t_k}+s b\left(Y_{t_{k}}, t_{k}\right)+\sqrt{s}\epsilon_{k+1},
~
k=0,1,\ldots, K-1,
\end{align}
where   $\{\epsilon_{k}\}_{k=1}^{K}$ are i.i.d. $N(0,\bI_{p})$ and
\begin{align}\label{driftb}
b(Y_{t_{k}},t_{k})=\frac{\Ebb_{Z}[\nabla f(Y_{t_{k}}+\sqrt{1-t_{k}}Z)]}{\Ebb_{Z}[f(Y_{t_{k}}+\sqrt{1-t_{k}} Z)]}=\frac{\Ebb_{Z}[Z f(Y_{t_{k}}+\sqrt{1-t_{k}}Z)]}{\Ebb_{Z}[f(Y_{t_{k}}+\sqrt{1-t_{k}} Z)]\sqrt{1-t_k}}.
\end{align}
From the definition of $b(Y_{t_{k}},t_{k})$  in (\ref{driftb})  we may not get its explicit expression.
Here, we can get one estimator $\tilde{b}_{m}$ of $b$ by replacing $\Ebb_{Z}$ in $b$ with $m$-sample mean, i.e.,
\begin{align}\label{drifte1}
\tilde{b}_m(Y_{t_{k}},t_{k})=
\frac{\frac{1}{m}\sum_{j=1}^m[\nabla f(Y_{t_{k}}+\sqrt{1-t_{k}}Z_j)]}{\frac{1}{m}\sum_{j=1}^{m} [f(Y_{t_{k}}+\sqrt{1-t_{k}} Z_j)]}, \ k=0, \ldots, K-1,
\end{align}
or
\begin{align}\label{drifte2}
\tilde{b}_m(Y_{t_{k}},t_{k})=
\frac{\frac{1}{m}\sum_{j=1}^m[Z_j f(Y_{t_{k}}+\sqrt{1-t_{k}}Z_j)]}{\frac{1}{m}\sum_{j=1}^{m} [f(Y_{t_{k}}+\sqrt{1-t_{k}} Z_j)]\cdot\sqrt{1-t_k}}, \ k=0, \ldots, K-1,
\end{align}
where  $Z_1, \ldots, Z_{m}$ are i.i.d. $N(0, \bI_p)$.
The detailed description of SFS is summarized in following Algorithm \ref{alg:1} below, which is  Algorithm 2 in \cite{sfs21}.
\begin{algorithm}[H]
	\caption{SFS for $\mu = \exp (-V(x))/C$  with  Monte Carlo estimation of the drift term}
    \label{alg:1}
	\begin{algorithmic}[1]
\STATE Input:  $m$, $K$.  Initialize $s=1/K$, $\widetilde{Y}_{t_0}=0$.
\FOR{$k= 0,1,\ldots, K-1$ }
\STATE Sample $\epsilon_{k+1}\sim N(0,\bI_{p})$.
\STATE Sample $Z_i, i,\ldots,m$, from $ N(0,\bI_{p})$.
\STATE Compute $\tilde{b}_{m}$ according to (\ref{drifte1}) or (\ref{drifte2}),
\STATE $
\widetilde{Y}_{t_{k+1}}=\widetilde{Y}_{t_{k}}+s\tilde{b}_{m}\left(\widetilde{Y}_{t_{k}}, t_{k}\right)+\sqrt{s} \epsilon_{k+1}$.
\ENDFOR
\STATE Output:  $\{\widetilde{Y}_{t_k}\}_{k=1}^{K}$.
\end{algorithmic}
\end{algorithm}

In Section 4.2 of \cite{sfs21}, they proved that
$$W_2(\mbox{Law}(\widetilde{Y}_{t_K}),\mu)\rightarrow 0, \ \ \mathrm{as} \ \  s \rightarrow 0, m \rightarrow \infty$$ under a restricted assumption that
the potential is uniformly strongly convex, i.e, $U(x,t)$ satisfies
(\ref{cond4}).
However, (\ref{cond4}) is not easy to verify.
In the next section, we establish a nonasymptotic bound on the  Wasserstein distance between  the  law of  $\widetilde{Y}_{t_K}$  generated by SFS (Algorithm \ref{alg:1}) and the target $\mu$ under smooth and bounded  conditions (\textbf{A1}) and (\textbf{A2}) but
 without using
{the strongly uniform convexity assumption (\ref{cond4}) on  $U(x,t)$.
}
\section{Nonasymptotic Bound  without convexity}\label{Theorey}
Under conditions (\textbf{A1}) and (\textbf{A2}), one can easily deduce the growth condition and Lipschitz/H{\"o}lder continuity of the drift term $b(x,t)$ \cite{sfs21}, i.e.,
\begin{align}\label{cond1}
\|b(x,t)\|_2^2\leq C_0(1+\|x\|_2^2), \tag{C1}
\end{align}
and
\begin{align}\label{cond2}
\|b(x,t)-b(y,t)\|_2\leq C_1 \|x-y\|_2 \tag{C2},
\end{align}
and
\begin{align}\label{cond3}
\|b(x,t)-b(y,s)\|_2\leq  C_1 \left(\|x-y\|_2+|t-s|^{\frac{1}{2}}\right), \tag{C3}
\end{align}
where $C_0$ and $C_1$ are two finite and positive constants.
\begin{remark}\label{R3}
(\ref{cond1}) and (\ref{cond2}) are the essentially sufficient
conditions such that  the
Schr{\"o}dinger-F\"{o}llmer SDE (\ref{sch-equation})
admits the unique strong solution.
 (\ref{cond3}) has  been introduced in  Theorem 4.1 of \cite{tzen2019theoretical}, and it is also similar to the condition H2 of \cite{chau2019stochastic} and Assumption 3.2 of \cite{barkhagen2018stochastic}.
 Obviously,  (\ref{cond3}) implies  (\ref{cond2}),  and (\ref{cond1}) holds if the drift term $b(x,t)$ is bounded over $\mathbb{R}^p\times[0,1]$. 
\end{remark}

  Let $\mathcal{D}(\nu_1, \nu_2)$ be the collection of coupling probability measures on $\left(\mathbb{R}^{2p},\mathcal{B}(\mathbb{R}^{2p})\right)$ such that its respective marginal distributions are $\nu_1$ and $\nu_2$.
The Wasserstein of order $d \geq 1$ with which we measure the discrepancy between $\mbox{Law}(\widetilde{Y}_{t_K})$  and $\mu$  is defined as
\begin{align*}
W_{d}(\nu_1, \nu_2)=\inf _{\nu \in \mathcal{D}(\nu_1, \nu_2)}\left(\int_{\mathbb{R}^{p}} \int_{\mathbb{R}^{p}}\left\|\theta_1-\theta_2\right\|_2^{d} \mathrm{d} \nu\left(\theta_1,\theta_2\right)\right)^{1/d}.
\end{align*}
\begin{theorem}\label{th1}
Assume (\textbf{A1}) and  (\textbf{A2}) hold, then
\begin{align*}
W_2(\mbox{Law}(\widetilde{Y}_{t_K}),\mu)\leq \mathcal{O}(\sqrt{ps})+\mathcal{O}\left(\sqrt{\frac{p}{\log(m)}}\right),
\end{align*}
where $s=1/K$ is the step size.
\end{theorem}

\begin{remark}
This theorem provides some guidance on the selection of $s$ and $m$. To ensure convergence of the distribution of $\widetilde{Y}_{t_K}$, we should set the step size $s = o(1/p)$ and
$m=\exp(p/o(1))$.
In high-dimensional models with a large $p$, we need to generate a large number of random vectors from $N(0, \bI_p)$ to obtain an accurate estimate of the drift term $b$.
If  we assume that $f$ is bounded above 
we can improve the nonasymptotic error bound, in which
$\mathcal{O}\left(\sqrt{p/\log(m)}\right)$
can be improved to be
$\mathcal{O}\left(\sqrt{p/m}\right)$.
\end{remark}
\begin{theorem}\label{th2}
Assume that, in addition to the conditions of Theorem \ref{th1}, $f$ has  a finite upper bound, then
\begin{align*}
W_2(\mbox{Law}(\widetilde{Y}_{t_K}),\mu)\leq \mathcal{O}(\sqrt{ps})+\mathcal{O}\left(\sqrt{\frac{p}{m}}\right),
\end{align*}
where $s=1/K$ is the step size.
\end{theorem}

\begin{remark}
With the boundedness condition on $f$, to ensure convergence of the sampling distribution, we can set the step size $s = o(1/p)$ and
$m=p/o(1)$.
Note that the sample size requirement for approximating the drift term is significantly less
stringent than that in Theorem \ref{th1}.
\end{remark}
\begin{remark}
Langevin sampling method has been studied under  the (strongly) convex potential assumption
\cite{durmus2016high-dimensional,durmus2016sampling,durmus2017nonasymptotic,dalalyan2017further,
dalalyan2017theoretical,cheng2018convergence,dalalyan2019user-friendly};
the dissipativity condition for the drift term \cite{raginsky2017non,mou2019improved,zhang2019nonasymptotic};
  the local convexity condition for the potential function outside a ball \cite{durmus2017nonasymptotic,cheng2018sharp,ma2019sampling,bou2020coupling}.
However, these conditions may not hold for models with multiple modes, for example,   Gaussian mixtures, where their potentials are not convex and the log Sobolev inequality may  not be satisfied.
Moreover, the constant in the log Sobolev inequality depends on the dimensionality exponentially \cite{wang2009log,hale2010asymptotic,menz2014poincare,raginsky2017non}, implying that the  Langevin samplers suffers from the curse of dimensionality.
 SFS does not require the underlying Markov process to be ergodic, therefore, our results in Theorem \ref{th1} and \ref{th2}
 are established under  the smooth and bounded  assumptions (\textbf{A1}) and (\textbf{A2}) on $f$
  but do not  need the above mentioned  conditions used in the analysis of Langevin samplers.
\end{remark}

In Theorem \ref{th1} and Theorem \ref{th2}, we use  (\textbf{A2}), i.e,  $f$ has positive lower bound,
however, (\textbf{A2}) may  not hold if the target distribution admits compact support. To circumvent this difficulty,  we consider the regularized probability  measure $$\mu_{\varepsilon}=(1-\varepsilon)\mu+\varepsilon G, \ \ \varepsilon\in(0,1).$$ The corresponding density ratio is
$$f_{\varepsilon} = \frac{d \mu_{\varepsilon}}{d G}=(1-\varepsilon)f+\varepsilon.$$ Obviously,
$f_{\varepsilon}$ satisfies (\textbf{A1}) and (\textbf{A2}) if   $f$ and $\nabla f$ are Lipschitz continuous.   Since $\mu_{\varepsilon}$ can approximate to $\mu$ well if we set $\varepsilon$ small enough, then
we consider sampling from $\mu_{\varepsilon}$ by running  SFS (Algorithm \ref{alg:1}) with $f$ being replaced by  $f_{\varepsilon}$.
We use $\widetilde{Y}_{t_K}^{\varepsilon}$ to denote  the last iteration of SFS.
\begin{theorem}\label{th3}
Assume (\textbf{A1}) holds and set $\varepsilon=(\log(m))^{-1/5}$, then
\begin{align*}
W_2(\mbox{Law}(\widetilde{Y}_{t_K}^{\varepsilon}),\mu)\leq \mathcal{O}(\sqrt{ps})+\widetilde{C}_p\cdot\mathcal{O}\left(\frac{1}{(\log(m))^{1/10}}\right),
\end{align*}
where  $s=1/K$ is the step size, $\widetilde{C}_p$ is a constant depending on $p$.
Moreover, if $f$ has the finite upper bound and set $\varepsilon=m^{-1/5}$, then
\begin{align*}
W_2(\mbox{Law}(\widetilde{Y}_{t_K}^{\varepsilon}),\mu)\leq \mathcal{O}(\sqrt{ps})+\widetilde{C}_p\cdot\mathcal{O}\left(\frac{1}{m^{1/10}}\right).
\end{align*}
\end{theorem}
\section{Conclusion}\label{conlusion}

In \cite{sfs21}, Schr\"{o}dinger-F\"{o}llmer sampler (SFS) was proposed  for sampling from possibly unnormalized distributions. The key feature of SFS is that it does not need ergodicity as its theoretical basis.
The consistency of SFS proved in \cite{sfs21} relies a restricted assumption that  
the potential function is uniformly  strongly convex.
In this paper we provide a new convergence  analysis of the SFS  without the  strongly  convexity condition on the potential.
 We establish a nonasymptotic error bound on Wasserstein distance between the law of the output of SFS and the target distribution
under   smooth and bounded assumptions on the   density ratio of the target distribution  over the standard normal distribution.

\section{Acknowledgment}
The authors would like to thank Professor Liming Wu at Universit\'{e} Clermont-Auvergne
for helpful discussions on this topic.

\begin{appendix}
\section{Appendix} \label{append}
\setcounter{equation}{0}
\def\theequation{A.\arabic{equation}}
In this appendix, we prove
Proposition  \ref{SBP} and Theorems \ref{th1}-\ref{th3}.

\subsection{Proof of Proposition  \ref{SBP}}
\begin{proof}
This is a known result, see for example \cite{dai1991stochastic,lehec2013representation}.
\end{proof}
\subsection{Preliminary lemmas for Theorems \ref{th1}-\ref{th2}}
First, we introduce  Lemmas \ref{lemma2}-\ref{lemma6} in preparing for the
proofs of Theorems \ref{th1}-\ref{th2}.
\begin{lemma}\label{lemma2}
Assume (\textbf{A1}) and (\textbf{A2}) hold,
then
\begin{align*}
\Ebb[\|X_t\|_2^2]\leq 2(C_0+p)\exp(2C_0t).
\end{align*}
\end{lemma}
\begin{proof}
From the definition of $X_t$ in (\ref{sch-equation}), we have
$
\|X_t\|_2\leq \int_{0}^t\|b(X_u,u)\|_2\mathrm{d}u+\|B_t\|_2.
$
Then, we can get
\begin{align*}
\|X_t\|_2^2&\leq
2\left(\int_{0}^t\|b(X_u,u)\|_2\mathrm{d}u\right)^2+2\|B_t\|_2^2\\
&\leq
2t\int_{0}^t\|b(X_u,u)\|_2^2\mathrm{d}u+2\|B_t\|_2^2\\
&\leq
2t\int_{0}^tC_0[\|X_u\|_2^2+1]\mathrm{d}u+2\|B_t\|_2^2,
\end{align*}
where the first inequality holds by the inequality $(a+b)^2\leq 2a^2+2b^2$, the last inequality holds by (\ref{cond1}).
Thus,
\begin{align*}
\Ebb\|X_t\|_2^2&\leq
2t\int_{0}^tC_0(\Ebb\|X_u\|_2^2+1)\mathrm{d}u+2\Ebb\|B_t\|_2^2\\
&\leq
2C_0\int_{0}^t \Ebb\|X_u\|_2^2\mathrm{d}u+2(C_0+p).
\end{align*}
By Bellman-Gronwall inequality, we have
\begin{align*}
\Ebb\|X_t\|_2^2\leq 2(C_0+p)\exp(2C_0t).
\end{align*}
\end{proof}
\begin{lemma}\label{lemma3}
Assume (\textbf{A1}) and (\textbf{A2}) hold,
then for any $0\leq t_1\leq t_2\leq 1$,
\begin{align*}
\Ebb[\|X_{t_2}-X_{t_1}\|_2^2]\leq 4C_0\exp(2C_0)(C_0+p)(t_2-t_1)^2+2C_0(t_2-t_1)^2+2p(t_2-t_1).
\end{align*}
\end{lemma}
\begin{proof}
From the definition of $X_t$ in (\ref{sch-equation}), we have
\begin{align*}
\|X_{t_2}-X_{t_1}\|_2\leq \int_{t_1}^{t_2}\|b(X_u,u)\|_2\mathrm{d}u+\|B_{t_2}-B_{t_1}\|_2.
\end{align*}
Then, we can get
\begin{align*}
\|X_{t_2}-X_{t_1}\|_2^2&\leq
2\left(\int_{t_1}^{t_2}\|b(X_u,u)\|_2\mathrm{d}u\right)^2+2\|B_{t_2}-B_{t_1}\|_2^2\\
&\leq
2(t_2-t_1)\int_{t_1}^{t_2}\|b(X_u,u)\|_2^2\mathrm{d}u+2\|B_{t_2}-B_{t_1}\|_2^2\\
&\leq
2(t_2-t_1)\int_{t_1}^{t_2}C_0[\|X_u\|_2^2+1]\mathrm{d}u+2\|B_{t_2}-B_{t_1}\|_2^2,
\end{align*}
where the last inequality holds by (\ref{cond1}).
Hence,
\begin{align*}
\Ebb\|X_{t_2}-X_{t_1}\|_2^2&\leq
2(t_2-t_1)\int_{t_1}^{t_2}C_0(\Ebb\|X_u\|_2^2+1)\mathrm{d}u+2\Ebb\|B_{t_2}-B_{t_1}\|_2^2\\
&\leq
4C_0\exp(2C_0)(C_0+p)(t_2-t_1)^2+2C_0(t_2-t_1)^2+2p(t_2-t_1),
\end{align*}
where the last inequality holds by Lemma \ref{lemma2}.
\end{proof}
\begin{lemma}\label{lemma4}
Assume (\textbf{A1}) and (\textbf{A2}) hold,
then for any $R>0$,
\begin{align*}
\underset{\|x\|_2
\leq R,t\in [0,1]}{\sup} \Ebb\left[\|b(x,t)-\tilde{b}_m(x,t)\|_2^2\right]
\leq \mathcal{O}\left(\frac{p\exp(R^2)}{m}\right).
\end{align*}
Moreover, if $f$ has the finite upper bound, then
\begin{align*}
\underset{x\in\mathbb{R}^p,t\in [0,1]}{\sup} \mathbb{E}\left[\|b(x,t)-\tilde{b}_m(t,x)\|_2^2\right]
\leq \mathcal{O}\left(\frac{p}{m}\right).
\end{align*}
\end{lemma}
\begin{proof}
Denote two independent sets of independent copies of $Z\sim N(0, \bI_p)$,
that is, $\bZ=\{Z_1,\ldots,Z_m\}$ and $\bZ^{\prime}=\{Z_1^{\prime},\ldots,Z_m^{\prime}\}$.
For notation convenience, we  denote
\begin{align*}
&d=\Ebb_{Z}\nabla f(x+\sqrt{1-t}Z),~ d_m=\frac{\sum_{i=1}^m\nabla f(x+\sqrt{1-t}Z_i)}{m},\\
&e=\Ebb_{Z} f(x+\sqrt{1-t}Z),~ e_m=\frac{\sum_{i=1}^m f(x+\sqrt{1-t}Z_i)}{m},\\
&d_m^{\prime}=\frac{\sum_{i=1}^m\nabla f(x+\sqrt{1-t}Z_i^{\prime})}{m},~e_m^{\prime}=\frac{\sum_{i=1}^m f(x+\sqrt{1-t}Z_i^{\prime})}{m}.
\end{align*}
Due to $d-d_m=\Ebb\left[d_m^{\prime}-d_m|\bZ\right]$,
then $\|d-d_m\|^2_2\leq \Ebb\left[\|d_m^{\prime}-d_m\|^2_2|\bZ\right]$.
Then,
\begin{align}\label{mm1}
\Ebb\|d-d_m\|^2
&\leq \Ebb\left[\Ebb[\|d_m^{\prime}-d_m\|^2_2|\bZ]\right]
=\Ebb\|d_m^{\prime}-d_m\|^2_2\notag\\
&=\frac{\Ebb_{Z_1,Z_1^{\prime}}\left\|\nabla f(x+\sqrt{1-t}Z_1)-\nabla f(x+\sqrt{1-t}Z_1^{\prime})\right\|^2_2}{m}\notag\\
&\leq \frac{(1-t)\gamma^2}{m}\Ebb_{Z_1,Z_1^{\prime}}\left\|Z_1-Z_1^{\prime}\right\|^2_2\notag\\
&\leq \frac{2p\gamma^2}{m},
\end{align}
where the second inequality holds by (\textbf{A1}). 
Similarly, we also have
\begin{align}\label{mm2}
\Ebb|e-e_m|^2
&\leq \Ebb|e_m^{\prime}-e_m|^2\notag\\
&=\frac{\Ebb_{Z_1,Z_1^{\prime}}
\left|f(x+\sqrt{1-t}Z_1)-f(x+\sqrt{1-t}Z_1^{\prime})\right|^2}{m}\notag\\
&\leq \frac{(1-t)\gamma^2}{m}\Ebb_{Z_1,Z_1^{\prime}}\left\|Z_1-Z_1^{\prime}\right\|^2_2\notag\\
&\leq \frac{2p\gamma^2}{m},
\end{align}
where the second inequality holds due to (\textbf{A1}). 
Thus, by (\ref{mm1}) and (\ref{mm2}), it follows that
\begin{align}\label{mm3}
\underset{x\in \mathbb{R}^p,t\in [0,1]}{\sup}\Ebb\left\|d-d_m\right\|_2^2
\leq \frac{2p\gamma^{2}}{m},
\end{align}
\begin{align}\label{mm4}
\underset{x\in \mathbb{R}^p,t\in [0,1]}{\sup}\Ebb|e-e_m|^2\leq \frac{2p\gamma^2}{m}.
\end{align}
Then, by (\textbf{A1}) and (\textbf{A2}),
through some simple calculation, it yields that
\begin{align}\label{mm5}
\|b(x,t)-\tilde{b}_m(x,t)\|_2
&=\left\|\frac{d}{e}-\frac{d_m}{e_m}\right\|_2\notag\\
&\leq \frac{\|d\|_2|e_m-e|+\|d-d_m\|_2|e|}{|ee_m|}\notag\\
&\leq \frac{\gamma|e_m-e|+\|d-d_m\|_2|e|}{\xi^2}.
\end{align}
Let $R>0$, then
\begin{align}\label{mm6}
\underset{\|x\|_2\leq R}\sup f(x)\leq \mathcal{O}\left(\exp(R^2/2)\right).
\end{align}
Therefore, by (\ref{mm3})-(\ref{mm6}), it can be concluded  that
\begin{align*}
\underset{\|x\|_2 \leq R,t\in [0,1]}{\sup} \Ebb\left[\|b(x,t)-\tilde{b}_m(x,t)\|_2^2\right]
\leq \mathcal{O}\left(\frac{p\exp(R^2)}{m}\right).
\end{align*}

Moreover, if $f$ has the finite upper bound, that is, there exists one finite and positive constant $\zeta$ such that $f\leq\zeta$. Then, similar to (\ref{mm5}), it follows that for all $x\in \mathbb{R}^p$ and $t\in[0,1]$,
\begin{align}\label{mm55}
\|b(x,t)-\tilde{b}_m(x,t)\|_2^2
&\leq  2\frac{\gamma^2|e_m-e|^2+\zeta^2\|d-d_m\|_2^2}{\xi^4}.
\end{align}
Then, by (\ref{mm3})-(\ref{mm4}) and (\ref{mm55}), it follows that
\begin{align*}
\underset{x\in\mathbb{R}^p,t\in [0,1]}{\sup} \mathbb{E}\left[\|b(x,t)-\tilde{b}_m(t,x)\|_2^2\right]
\leq \mathcal{O}\left(\frac{p}{m}\right).
\end{align*}
\end{proof}
\begin{lemma}\label{lemma5}
Assume (\textbf{A1}) and (\textbf{A2}) hold,
then for $k=0,1,\ldots,K$,
\begin{align*}
E[\|\widetilde{Y}_{t_{k}}\|^2_2]
\leq  \frac{6\gamma^2}{\xi^2}+3p.
\end{align*}
\end{lemma}
\begin{proof}
Define
$\Theta_{k,t}=\widetilde{Y}_{t_k}+(t-t_k)\tilde{b}_m(\widetilde{Y}_{t_k},t_k)$
and $\widetilde{Y}_{t}=\Theta_{k,t}+B_t-B_{t_k}$, where $t_k \leq t \leq t_{k+1}$ with $k=0,1,\ldots,K-1$.
By (\textbf{A1}) and (\textbf{A2}), it follows that for all $x \in \mathbb{R}^p$ and $t\in[0,1]$,
\begin{align}\label{eq2}
\|b(x,t)\|_2^2\leq \frac{\gamma^2}{\xi^2},~~ \|\tilde{b}_m(x,t)\|_2^2 \leq \frac{\gamma^2}{\xi^2}.
\end{align}
Then, by (\ref{eq2}),  we have
\begin{align*}
\|\Theta_{k,t}\|^2_2&=\|\widetilde{Y}_{t_k}\|^2_2+(t-t_k)^2\|\tilde{b}_m(\widetilde{Y}_{t_k},t_k)\|^2_2
+2(t-t_k)\widetilde{Y}_{t_k}^{\top}\tilde{b}_m(\widetilde{Y}_{t_k},t_k)\\
&\leq (1+s)\|\widetilde{Y}_{t_k}\|^2_2+\frac{(s+s^2)\gamma^2}{\xi^2}.
\end{align*}
Further, we can get
\begin{align*}
\Ebb[\|\widetilde{Y}_{t}\|^2_2|\widetilde{Y}_{t_k}]&=\Ebb[\|\Theta_{k,t}\|^2_2|\widetilde{Y}_{t_k}]+(t-t_k)p\\
&\leq(1+s)\Ebb\|\widetilde{Y}_{t_k}\|^2_2+\frac{(s+s^2)\gamma^2}{\xi^2}+sp.
\end{align*}
Therefore,
\begin{align*}
\Ebb[\|\widetilde{Y}_{t_{k+1}}\|^2_2]
\leq(1+s)\Ebb\|\widetilde{Y}_{t_k}\|^2_2+\frac{(s+s^2)\gamma^2}{\xi^2}+sp.
\end{align*}
Since $\widetilde{Y}_{t_0}=0$, then by induction, we have
\begin{align*}
\Ebb[\|\widetilde{Y}_{t_{k+1}}\|^2_2]
\leq \frac{6\gamma^2}{\xi^2}+3p.
\end{align*}
\end{proof}
\begin{lemma}\label{lemma6}
Assume (\textbf{A1}) and (\textbf{A2}) hold,
then for $k=0,1,\ldots,K$,
\begin{align*}
\Ebb\left\|b(\widetilde{Y}_{t_k},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k},t_k)\right\|_2^2\leq
\mathcal{O}\left(\frac{p}{\log(m)}\right).
\end{align*}
Moreover, if $f$  has the finite upper bound, then 
\begin{align*}
\Ebb\left\|b(\widetilde{Y}_{t_k},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k},t_k)\right\|_2^2\leq
\mathcal{O}\left(\frac{p}{m}\right).
\end{align*}
\end{lemma}
\begin{proof}
Let $R>0$, then
\begin{equation}\label{erro1}
\begin{split}
\Ebb\left\|b(\widetilde{Y}_{t_k},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k},t_k)\right\|_2^2
&=\Ebb_{\widetilde{Y}_{t_k}}\Ebb_Z\left[\left\|b(\widetilde{Y}_{t_k},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k},t_k)\right\|_2^21(\|\widetilde{Y}_{t_k}\|_2\leq R)\right]\\
&~~~+\Ebb_{\widetilde{Y}_{t_k}}\Ebb_Z\left[\left\|b(\widetilde{Y}_{t_k},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k},t_k)\right\|_2^21(\|\widetilde{Y}_{t_k}\|_2> R)\right].
\end{split}
\end{equation}
Next, we need to bound the two terms of (\ref{erro1}).
First, by Lemma \ref{lemma4}, we have
$$
\Ebb_{\widetilde{Y}_{t_k}}\Ebb_Z\left[\left\|b(\widetilde{Y}_{t_k},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k},t_k)\right\|_2^21(\|\widetilde{Y}_{t_k}\|_2\leq R)\right]\leq  \mathcal{O}\left(\frac{p\exp(R^2)}{m}\right).
$$
Secondly, combining (\ref{eq2}) and Lemma \ref{lemma5} with Markov inequality, it yields that
$$
\Ebb_{\widetilde{Y}_{t_k}}\Ebb_Z\left[\left\|b(\widetilde{Y}_{t_k},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k},t_k)\right\|_2^21(\|\widetilde{Y}_{t_k}\|_2> R)\right]
\leq \mathcal{O}\left(p/R^2\right).
$$
Thence
\begin{align}\label{err1}
\Ebb\left\|b(\widetilde{Y}_{t_k},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k},t_k)\right\|_2^2
\leq \mathcal{O}\left(\frac{p\exp(R^2)}{m}\right)
+\mathcal{O}\left(p/R^2\right).
\end{align}
Set $R=\left(\frac{\log(m)}{2}\right)^{1/2}$ in (\ref{err1}), then we have
\begin{align*}
\Ebb\left\|b(\widetilde{Y}_{t_k},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k},t_k)\right\|_2^2\leq
\mathcal{O}\left(\frac{p}{\log(m)}\right).
\end{align*}

Moreover, if $f$ has the finite upper bound, then by Lemma \ref{lemma4}, we can similarly get
$$
\Ebb\left\|b(\widetilde{Y}_{t_k},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k},t_k)\right\|_2^2=
\Ebb_{\widetilde{Y}_{t_k}}\Ebb_{Z}\left[\left\|b(\widetilde{Y}_{t_k},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k},t_k)\right\|_2^2\right]
\leq \mathcal{O}\left(\frac{p}{m}\right).
$$
This completes the proof.
\end{proof}
\subsection{Proof of Theorem \ref{th1}}
\begin{proof}
From the definition of $\widetilde{Y}_{t_k}$ and $X_{t_k}$, we have
\begin{align*}
&\|\widetilde{Y}_{t_k}-X_{t_k}\|_2^2\\
&\leq \|\widetilde{Y}_{t_{k-1}}-X_{t_{k-1}}\|_2^2
+\left(\int_{t_{k-1}}^{t_k}\|b(X_u,u)-\tilde{b}_m(\widetilde{Y}_{t_{k-1}},t_{k-1})\|_2\mathrm{d} u\right)^2\\
&~~~+2\|\widetilde{Y}_{t_{k-1}}-X_{t_{k-1}}\|_2
\left(\int_{t_{k-1}}^{t_k}\|b(X_u,u)-\tilde{b}_m(\widetilde{Y}_{t_{k-1}},t_{k-1})\|_2\mathrm{d} u\right)\\
&\leq (1+s) \|\widetilde{Y}_{t_{k-1}}-X_{t_{k-1}}\|_2^2
+(1+s)\int_{t_{k-1}}^{t_k}\|b(X_u,u)-\tilde{b}_m(\widetilde{Y}_{t_{k-1}},t_{k-1})\|_2^2\mathrm{d} u\\
&\leq (1+s) \|\widetilde{Y}_{t_{k-1}}-X_{t_{k-1}}\|_2^2
+2(1+s)\int_{t_{k-1}}^{t_k}\|b(X_u,u)-b(\widetilde{Y}_{t_{k-1}},t_{k-1})\|_2^2\mathrm{d} u\\
&~~~+2s(1+s)\|b(\widetilde{Y}_{t_{k-1}},t_{k-1})-\tilde{b}_m(\widetilde{Y}_{t_{k-1}},t_{k-1})\|_2^2\\
&\leq (1+s) \|\widetilde{Y}_{t_{k-1}}-X_{t_{k-1}}\|_2^2+4C_1^2(1+s)\int_{t_{k-1}}^{t_k}[\|X_u-\widetilde{Y}_{t_{k-1}}\|_2^2+|u-t_{k-1}|]\mathrm{d} u\\
&~~~+2s(1+s)\|b(\widetilde{Y}_{t_{k-1}},t_{k-1})-\tilde{b}_m(\widetilde{Y}_{t_{k-1}},t_{k-1})\|_2^2\\
&\leq (1+s) \|\widetilde{Y}_{t_{k-1}}-X_{t_{k-1}}\|_2^2
+8C_1^2(1+s)\int_{t_{k-1}}^{t_k}\|X_u-X_{t_{k-1}}\|_2^2\mathrm{d} u\\
&~~~+8C_1^2s(1+s)\|X_{t_{k-1}}-\widetilde{Y}_{t_{k-1}}\|_2^2+4C_1^2(1+s)s^2\\
&~~~+2s(1+s)\|b(\widetilde{Y}_{t_{k-1}},t_{k-1})-\tilde{b}_m(\widetilde{Y}_{t_{k-1}},t_{k-1})\|_2^2\\
&\leq (1+s+8C_1^2(s+s^2))\|\widetilde{Y}_{t_{k-1}}-X_{t_{k-1}}\|_2^2
+8C_1^2(1+s)\int_{t_{k-1}}^{t_k}\|X_u-X_{t_{k-1}}\|_2^2\mathrm{d} u\\
&~~~+4C_1^2(1+s)s^2+2s(1+s)\|b(\widetilde{Y}_{t_{k-1}},t_{k-1})-\tilde{b}_m(\widetilde{Y}_{t_{k-1}},t_{k-1})\|_2^2,
\end{align*}
where the second inequality holds duo to $2ab\leq s a^2+\frac{b^2}{s}$,
the fourth inequality holds by (\ref{cond3}).
Then,
\begin{align}\label{th1eq1}
\Ebb\|\widetilde{Y}_{t_k}-X_{t_k}\|_2^2
&\leq (1+s+8C_1^2(s+s^2)) \Ebb\|\widetilde{Y}_{t_{k-1}}-X_{t_{k-1}}\|_2^2\notag\\
&~~~+8C_1^2(1+s)\int_{t_{k-1}}^{t_k}\Ebb\|X_u-X_{t_{k-1}}\|_2^2\mathrm{d} u+4C_1^2(s^2+s^3)\notag\\
&~~~+2s(1+s)\Ebb[\|b(\widetilde{Y}_{t_{k-1}},t_{k-1})-\tilde{b}_m(\widetilde{Y}_{t_{k-1}},t_{k-1})\|_2^2]\notag\\
&\leq(1+s+8C_1^2(s+s^2)) \Ebb\|\widetilde{Y}_{t_{k-1}}-X_{t_{k-1}}\|_2^2+h(s)\notag\\
&~~~+4C_1^2(s^2+s^3)+2s(1+s)\Ebb[\|b(\widetilde{Y}_{t_{k-1}},t_{k-1})-\tilde{b}_m(\widetilde{Y}_{t_{k-1}},t_{k-1})\|_2^2]\notag\\
&\leq(1+s+8C_1^2(s+s^2)) \Ebb\|\widetilde{Y}_{t_{k-1}}-X_{t_{k-1}}\|_2^2+h(s)\notag\\
&~~~+4C_1^2(s^2+s^3)+2s(1+s)\mathcal{O}\left(\frac{p}{\log(m)}\right),
\end{align}
where $h(s)=8C_1^2(s+s^2)[4C_0\exp(2C_0)(C_0+p)s^2+2C_0s^2+2ps]$, and the last inequality holds by Lemma \ref{lemma6}.
Owing to $\widetilde{Y}_{t_0}=X_{t_0}=0$, we can conclude that
\begin{align*}
&\Ebb\|\widetilde{Y}_{t_K}-X_{t_K}\|_2^2\\
&\leq\frac{(1+s+8C_1^2(s+s^2))^K-1}{s+8C_1^2(s+s^2)}
\left[h(s)+4C_1^2(s^2+s^3)+2(s+s^2)\mathcal{O}\left(\frac{p}{\log(m)}\right)\right]\\
&\leq \mathcal{O}(ps)+\mathcal{O}\left(\frac{p}{\log(m)}\right).
\end{align*}
Therefore,
\begin{align*}
W_2(Law(\widetilde{Y}_{t_K}),\mu)
\leq \mathcal{O}(\sqrt{ps})+\mathcal{O}\left(\sqrt{\frac{p}{\log(m)}}\right).
\end{align*}
\end{proof}
\subsection{Proof of Theorem \ref{th2}}
\begin{proof}
This proof is same as that of Theorem \ref{th1}.  Similar to (\ref{th1eq1}), by Lemma \ref{lemma6}, it yields that
\begin{align*}
\Ebb\|\widetilde{Y}_{t_k}-X_{t_k}\|_2^2
&\leq(1+s+8C_1^2(s+s^2)) \Ebb\|\widetilde{Y}_{t_{k-1}}-X_{t_{k-1}}\|_2^2+h(s)\\
&~~~+4C_1^2(s^2+s^3)+2s(1+s)\mathcal{O}\left(\frac{p}{m}\right).
\end{align*}
Then,  we also have
\begin{align*}
&\Ebb\|\widetilde{Y}_{t_K}-X_{t_K}\|_2^2\\
&\leq\frac{(1+s+8C_1^2(s+s^2))^K-1}{s+8C_1^2(s+s^2)}
\left[h(s)+4C_1^2(s^2+s^3)+2(s+s^2)\mathcal{O}\left(\frac{1}{m}\right)\right]\\
&\leq \mathcal{O}(ps)+\mathcal{O}\left(\frac{p}{m}\right).
\end{align*}
Hence, it follows that
\begin{align*}
W_2(Law(\widetilde{Y}_{t_K}),\mu)
\leq \mathcal{O}(\sqrt{ps})+\mathcal{O}\left(\sqrt{\frac{p}{m}}\right).
\end{align*}
\end{proof}

\subsection{Preliminary lemmas for Theorem \ref{th3}}
To prove Theorem \ref{th3}, we first prove the Lemmas \ref{lemma41}-\ref{lemma61}.
\begin{lemma}\label{lemma41}
Assume (\textbf{A1}) holds,
then for any $R>0$,
\begin{align*}
\underset{\|x\|_2 \leq R,t\in [0,1]}{\sup} \Ebb\left[\|b(x,t)-\tilde{b}_m(x,t)\|_2^2\right]
\leq
\mathcal{O}\left(\frac{p\exp(R^2)(C_p)^4}{m\varepsilon^4}\right)
+\mathcal{O}\left(\frac{p(C_p)^2}{m\varepsilon^2}\right),
\end{align*}
where $C_p=(2\pi)^{p/2}C^{-1}$.
Moreover, if $f$ has the finite upper bound, then
\begin{align*}
\underset{x\in\mathbb{R}^p,t\in [0,1]}{\sup} \Ebb\left[\|b(x,t)-\tilde{b}_m(t,x)\|_2^2\right]
\leq \mathcal{O}\left(\frac{p(C_p)^4}{m\varepsilon^4}\right)
+\mathcal{O}\left(\frac{p(C_p)^2}{m\varepsilon^2}\right).
\end{align*}
\end{lemma}
\begin{proof}
Denote two independent sets of independent copies of $Z\sim N(0, \bI_p)$ by $\bZ=\{Z_1,\ldots,Z_m\}$ and $\bZ^{\prime}=\{Z_1^{\prime},\ldots,Z_m^{\prime}\}$.
For notation convenience, we  denote
\begin{align*}
&d=\Ebb_{Z}\nabla g(x+\sqrt{1-t}Z),~ d_m=\frac{\sum_{i=1}^m\nabla g(x+\sqrt{1-t}Z_i)}{m},\\
&e=\Ebb_{Z} \left[g(x+\sqrt{1-t}Z)+\frac{\varepsilon }{ C_p(1-\varepsilon)}\right],~ e_m=\frac{\sum_{i=1}^m g(x+\sqrt{1-t}Z_i)}{m}+\frac{\varepsilon}{ C_p(1-\varepsilon)},\\
&d_m^{\prime}=\frac{\sum_{i=1}^m\nabla g(x+\sqrt{1-t}Z_i^{\prime})}{m},~e_m^{\prime}=\frac{\sum_{i=1}^m g(x+\sqrt{1-t}Z_i^{\prime})}{m}+\frac{\varepsilon}{ C_p(1-\varepsilon)},
\end{align*}
where $g(x)=\exp(\|x\|^2_2/2-V(x))$.
Since  $d-d_m=\Ebb\left[d_m^{\prime}-d_m|\bZ\right]$,
we have  $\|d-d_m\|^2_2\leq \Ebb\left[\|d_m^{\prime}-d_m\|^2_2|\bZ\right]$.
By (\textbf{A1}),
it yields that $g$ and $\nabla g$ are Lipschitz continuous.
Thus  there exists a finite and positive constant $\gamma$
such that for all $x,y \in \mathbb{R}^p$,
\begin{align}\label{L11}
|g(x)-g(y)|\leq \gamma \|x-y\|_2,
\end{align}
\begin{align}\label{L21}
\|\nabla g(x)- \nabla g(y)\|_2\leq \gamma \|x-y\|_2.
\end{align}
Therefore,
\begin{align}\label{mm11}
\Ebb\|d-d_m\|^2
&\leq \Ebb\left[\Ebb[\|d_m^{\prime}-d_m\|^2_2|\bZ]\right]=\Ebb\|d_m^{\prime}-d_m\|^2_2\notag\\
&=\frac{\Ebb_{Z_1,Z_1^{\prime}}\left\|\nabla g(x+\sqrt{1-t}Z_1)-\nabla g(x+\sqrt{1-t}Z_1^{\prime})\right\|^2_2}{m}\notag\\
&\leq \frac{(1-t)\gamma^2}{m}\Ebb_{Z_1,Z_1^{\prime}}\left\|Z_1-Z_1^{\prime}\right\|^2_2\notag\\
&\leq \frac{2p\gamma^2}{m},
\end{align}
where the second inequality follows from (\ref{L21}).
Similarly, we also have
\begin{align}\label{mm21}
\Ebb|e-e_m|^2
&\leq \Ebb|e_m^{\prime}-e_m|^2\notag\\
&=\frac{\Ebb_{Z_1,Z_1^{\prime}}
\left|g(x+\sqrt{1-t}Z_1)-g(x+\sqrt{1-t}Z_1^{\prime})\right|^2}{m}\notag\\
&\leq \frac{(1-t)\gamma^2}{m}\Ebb_{Z_1,Z_1^{\prime}}\left\|Z_1-Z_1^{\prime}\right\|^2_2\notag\\
&\leq \frac{2p\gamma^2}{m},
\end{align}
where the second inequality follows from (\ref{L11}).
Hence, by (\ref{mm11}) and (\ref{mm21}), we have
\begin{align}\label{mm31}
\underset{x\in \mathbb{R}^p,t\in [0,1]}{\sup}\Ebb\left\|d-d_m\right\|_2^2
\leq \frac{2p\gamma^{2}}{m},
\end{align}
\begin{align}\label{mm41}
\underset{x\in \mathbb{R}^p,t\in [0,1]}{\sup}\Ebb|e-e_m|^2\leq \frac{2p\gamma^2}{m}.
\end{align}
Then, by (\ref{L11}) and (\ref{L21}), 
through some simple calculation, it yields that
\begin{align}\label{mm51}
\|b(x,t)-\tilde{b}_m(x,t)\|_2
&=\left\|\frac{d}{e}-\frac{d_m}{e_m}\right\|_2\notag\\
&\leq \frac{\|d\|_2|e_m-e|+\|d-d_m\|_2|e|}{|ee_m|}\notag\\
&\leq \frac{\gamma|e_m-e|+\|d-d_m\|_2|e|}{(\varepsilon /( C_p-C_p\varepsilon))^2}.
\end{align}
Let $R>0$, then
\begin{align}\label{mm61}
\underset{\|x\|_2\leq R}\sup g(x)\leq \mathcal{O}\left(\exp(R^2/2)\right).
\end{align}
Therefore, by (\ref{mm31})-(\ref{mm61}), it can be concluded  that
\begin{align*}
\underset{\|x\|_2 \leq R,t\in [0,1]}{\sup} \Ebb\left[\|b(x,t)-\tilde{b}_m(x,t)\|_2^2\right]
&\leq
\mathcal{O}\left(\frac{p\exp(R^2)}{m(\varepsilon /(C_p-C_p\varepsilon))^4}\right)
+
\mathcal{O}\left(\frac{p}{m(\varepsilon /(C_p-C_p\varepsilon))^2}\right)\\
&\leq
\mathcal{O}\left(\frac{p\exp(R^2)(C_p)^4}{m\varepsilon^4}\right)
+
\mathcal{O}\left(\frac{p(C_p)^2}{m\varepsilon^2}\right).
\end{align*}

Moreover,  $f$ has a finite upper bound so does $g$. Then there exists a finite and positive constant $\zeta$ such that $g\leq\zeta$. Similar to (\ref{mm51}), it follows that for all $x\in \mathbb{R}^p$ and $t\in[0,1]$,
\begin{align}\label{mm551}
\|b(x,t)-\tilde{b}_m(x,t)\|_2^2
&\leq  2\frac{\gamma^2|e_m-e|^2+(\zeta+\varepsilon /(C_p-C_p\varepsilon))^2\|d-d_m\|_2^2}{(\varepsilon /(C_p-C_p\varepsilon))^4}.
\end{align}
Then, by (\ref{mm31})-(\ref{mm41}) and (\ref{mm551}), it follows that
\begin{align*}
\underset{x\in\mathbb{R}^p,t\in [0,1]}{\sup} \Ebb\left[\|b(x,t)-\tilde{b}_m(t,x)\|_2^2\right]
&\leq \mathcal{O}\left(\frac{p}{m(\varepsilon /(C_p-C_p\varepsilon))^4}\right)+\mathcal{O}\left(\frac{p}{m(\varepsilon /(C_p-C_p\varepsilon))^2}\right)\\
&\leq \mathcal{O}\left(\frac{p(C_p)^4}{m\varepsilon^4}\right)+\mathcal{O}\left(\frac{p(C_p)^2}{m\varepsilon^2}\right).
\end{align*}
\end{proof}
\begin{lemma}\label{lemma51}
Assume (\textbf{A1}) holds,
then for $k=0,1,\ldots,K$,
\begin{align*}
\Ebb[\|\widetilde{Y}_{t_{k}}^{\varepsilon}\|^2_2]
\leq\mathcal{O}\left(\frac{(C_p)^2}{\varepsilon^2}\right)+\mathcal{O}\left(p\right),
\end{align*}
where $C_p=(2\pi)^{p/2}C^{-1}$.
\end{lemma}
\begin{proof}
Define
$\Theta_{k,t}^{\varepsilon}=\widetilde{Y}_{t_k}^{\varepsilon}+(t-t_k)\tilde{b}_m(\widetilde{Y}_{t_k}^{\varepsilon},t_k)$
and $\widetilde{Y}_{t}^{\varepsilon}=\Theta_{k,t}^{\varepsilon}+B_t-B_{t_k}$, where $t_k \leq t \leq t_{k+1}$ with $k=0,1,\ldots,K-1$.
By (\textbf{A1}), then there exists one finite and positive constant $\gamma$ such that  $g$ is $\gamma$-Lipschitz continuous.
Then, for all $x \in \mathbb{R}^p$ and $t\in[0,1]$, we have
\begin{align}\label{eq21}
\|b(x,t)\|_2^2\leq \frac{\gamma^2}{(\varepsilon /(C_p-C_p\varepsilon))^2
},~~ \|\tilde{b}_m(x,t)\|_2^2 \leq \frac{\gamma^2}{(\varepsilon /(C_p-C_p\varepsilon))^2
}.
\end{align}
By (\ref{eq21}),  we have
\begin{align*}
\|\Theta_{k,t}^{\varepsilon}\|^2_2&=\|\widetilde{Y}_{t_k}^{\varepsilon}\|^2_2+(t-t_k)^2\|\tilde{b}_m(\widetilde{Y}_{t_k}^{\varepsilon},t_k)\|^2_2
+2(t-t_k)(\widetilde{Y}_{t_k}^{\varepsilon})^{\top}\tilde{b}_m(\widetilde{Y}_{t_k}^{\varepsilon},t_k)\\
&\leq (1+s)\|\widetilde{Y}_{t_k}^{\varepsilon}\|^2_2+\frac{(s+s^2)\gamma^2}{(\varepsilon /(C_p-C_p\varepsilon))^2}.
\end{align*}
Furthermore, it can be shown that
\begin{align*}
\Ebb[\|\widetilde{Y}_{t}^{\varepsilon}\|^2_2|\widetilde{Y}_{t_k}^{\varepsilon}]
&=\Ebb[\|\Theta_{k,t}^{\varepsilon}\|^2_2|\widetilde{Y}_{t_k}^{\varepsilon}]+(t-t_k)p\\
&\leq(1+s)\Ebb\|\widetilde{Y}_{t_k}^{\varepsilon}\|^2_2+\frac{(s+s^2)\gamma^2}{(\varepsilon /(C_p-C_p\varepsilon))^2}+sp.
\end{align*}
Therefore,
\begin{align*}
\Ebb[\|\widetilde{Y}_{t_{k+1}}^{\varepsilon}\|^2_2]
\leq(1+s)\Ebb[\|\widetilde{Y}_{t_k}^{\varepsilon}\|^2_2]+\frac{(s+s^2)\gamma^2}{(\varepsilon /(C_p-C_p\varepsilon))^2}+sp.
\end{align*}
Since $\widetilde{Y}_{t_0}^{\varepsilon}=0$, then by induction, we have
\begin{align*}
\Ebb[\|\widetilde{Y}_{t_{k+1}}^{\varepsilon}\|^2_2]
\leq \frac{6\gamma^2}{(\varepsilon /(C_p-C_p\varepsilon))^2}+3p
\leq\mathcal{O}\left(\frac{(C_p)^2}{\varepsilon^2}\right)+\mathcal{O}\left(p\right).
\end{align*}
\end{proof}
\begin{lemma}\label{lemma61}
Assume (\textbf{A1})  holds,
then for $k=0,1,\ldots,K$ and $t\in [0,1]$,
\begin{align*}
\Ebb\left\|b(\widetilde{Y}_{t_k}^{\varepsilon},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k}^{\varepsilon},t_k)\right\|_2^2
&\leq
\mathcal{O}\left(\frac{p(C_p)^4}{\sqrt{m}\varepsilon^4}\right)
+\mathcal{O}\left(\frac{(C_p)^4}{\log(m)\varepsilon^4}\right)
+\mathcal{O}\left(\frac{p(C_p)^2}{\log(m)\varepsilon^2}\right),
\end{align*}
where $C_p=(2\pi)^{p/2}C^{-1}$.
Moreover, if $f$  has a finite upper bound, then
\begin{align*}
\Ebb\left\|b(\widetilde{Y}_{t_k}^{\varepsilon},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k}^{\varepsilon},t_k)\right\|_2^2
 \leq \mathcal{O}\left(\frac{p(C_p)^4}{m\varepsilon^4}\right)
+\mathcal{O}\left(\frac{p(C_p)^2}{m\varepsilon^2}\right).
\end{align*}
\end{lemma}
\begin{proof}
Let $R>0$, then
\begin{equation}\label{erro11}
\begin{split}
\Ebb\left\|b(\widetilde{Y}_{t_k}^{\varepsilon},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k}^{\varepsilon},t_k)\right\|_2^2
&=\Ebb_{\widetilde{Y}_{t_k}^{\varepsilon}}\Ebb_Z\left[\left\|b(\widetilde{Y}_{t_k}^{\varepsilon},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k}^{\varepsilon},t_k)\right\|_2^21(\|\widetilde{Y}_{t_k}^{\varepsilon}\|_2\leq R)\right]\\
&~~~+\Ebb_{\widetilde{Y}_{t_k}^{\varepsilon}}\Ebb_Z\left[\left\|b(\widetilde{Y}_{t_k}^{\varepsilon},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k}^{\varepsilon},t_k)\right\|_2^21(\|\widetilde{Y}_{t_k}^{\varepsilon}\|_2> R)\right].
\end{split}
\end{equation}
Next, we need to bound the two terms on the right hand of (\ref{erro11}).
First, by Lemma \ref{lemma41}, we have
\begin{align*}
\Ebb_{\widetilde{Y}_{t_k}^{\varepsilon}}\Ebb_Z\left[\left\|b(\widetilde{Y}_{t_k}^{\varepsilon},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k}^{\varepsilon},t_k)\right\|_2^21(\|\widetilde{Y}_{t_k}^{\varepsilon}\|_2\leq R)\right]
\leq
\mathcal{O}\left(\frac{p\exp(R^2)(C_p)^4}{m\varepsilon^4}\right)
+\mathcal{O}\left(\frac{p(C_p)^2}{m\varepsilon^2}\right).
\end{align*}
Second, by combining (\ref{eq21}) and Lemma \ref{lemma51} with the Markov inequality, we have
\begin{align*}
\Ebb_{\widetilde{Y}_{t_k}^{\varepsilon}}\Ebb_Z\left[\left\|b(\widetilde{Y}_{t_k}^{\varepsilon},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k}^{\varepsilon},t_k)\right\|_2^21(\|\widetilde{Y}_{t_k}^{\varepsilon}\|_2> R)\right]
&\leq \mathcal{O}\left(\frac{(C_p)^4}{R^2\varepsilon^4}\right)+\mathcal{O}\left(\frac{p(C_p)^2}{R^2\varepsilon^2}\right).
\end{align*}
Therefore,
\begin{align}\label{err11}
\Ebb\left\|b(\widetilde{Y}_{t_k}^{\varepsilon},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k}^{\varepsilon},t_k)\right\|_2^2
&\leq
\mathcal{O}\left(\frac{p\exp(R^2)(C_p)^4}{m\varepsilon^4}\right)
+\mathcal{O}\left(\frac{p(C_p)^2}{m\varepsilon^2}\right)\notag\\
&~~~+\mathcal{O}\left(\frac{(C_p)^4}{R^2\varepsilon^4}\right)
+\mathcal{O}\left(\frac{p(C_p)^2}{R^2\varepsilon^2}\right).
\end{align}
Setting $R=\left(\frac{\log(m)}{2}\right)^{1/2}$ in (\ref{err11}), we have
\begin{align*}
\Ebb\left\|b(\widetilde{Y}_{t_k}^{\varepsilon},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k}^{\varepsilon},t_k)\right\|_2^2
&\leq
\mathcal{O}\left(\frac{p(C_p)^4}{\sqrt{m}\varepsilon^4}\right)
+\mathcal{O}\left(\frac{(C_p)^4}{\log(m)\varepsilon^4}\right)
+\mathcal{O}\left(\frac{p(C_p)^2}{\log(m)\varepsilon^2}\right).
\end{align*}
Moreover, if $f$ has a finite upper bound, then by Lemma \ref{lemma41}, we can similarly get
\begin{align*}
\Ebb\left\|b(\widetilde{Y}_{t_k}^{\varepsilon},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k}^{\varepsilon},t_k)\right\|_2^2
&=\Ebb_{\widetilde{Y}_{t_k}^{\varepsilon}}\Ebb_{Z}\left[\left\|b(\widetilde{Y}_{t_k}^{\varepsilon},t_k)-\tilde{b}_m(\widetilde{Y}_{t_k}^{\varepsilon},t_k)\right\|_2^2\right]\\
& \leq \mathcal{O}\left(\frac{p(C_p)^4}{m\varepsilon^4}\right)
+\mathcal{O}\left(\frac{p(C_p)^2}{m\varepsilon^2}\right).
\end{align*}
This completes the proof.
\end{proof}
\subsection{Proof of Theorem \ref{th3}}
\begin{proof}
By triangle inequality, we have  $W_2(Law(\widetilde{Y}_{t_K}^{\varepsilon}),\mu)\leq
W_2(Law(\widetilde{Y}_{t_K}^{\varepsilon}),\mu_{\varepsilon})
+W_2(\mu,\mu_{\varepsilon})$,
then we  obtain the upper bound of two terms on the right hand of this inequality, respectively.

First,
similar to the proof of Theorem \ref{th1}, by Lemma \ref{lemma61} and $\widetilde{Y}_{t_0}^{\varepsilon}=X_{t_0}=0$ and through some calculation, we can conclude that
\begin{align}\label{ERRR1}
W_2(Law(\widetilde{Y}_{t_K}^{\varepsilon}),\mu_{\varepsilon})
\leq
\mathcal{O}(\sqrt{ps})+\mathcal{O}\left(\frac{\sqrt{p}(C_p)^2}{m^{1/4}\varepsilon^2}\right)
+\mathcal{O}\left(\frac{(C_p)^2}{\sqrt{\log(m)}\varepsilon^2}\right)
+\mathcal{O}\left(\frac{\sqrt{p}C_p}{\sqrt{\log(m)}\varepsilon}\right).
\end{align}

Second, we need to get the upper bound of $W_2(\mu,\mu_{\varepsilon})$.
Let $Y\sim \mu$ and $Z\sim N(0,\bI_p)$,  $\theta$ is one Bernoulli random variable satisfying  $P(\theta=1)=1-\varepsilon$ and
$P(\theta=0)=\varepsilon$. Assume $Y$, $Z$ and $\theta$ are independent of each other.
Then $(Y,(1-\theta)Z+\theta Y)$ is one coupling of $(\mu,\mu_{\varepsilon})$, and denote  its joint distribution by $\pi$. Therefore, we have
\begin{align*}
\int_{\mathbb{R}^p\times\mathbb{R}^p}\|x-y\|_2^2d \pi&=\Ebb\left\|Y-((1-\theta)Z+\theta Y)\right\|_2^2\\
&=\Ebb\left[\Ebb\left[\|Y-((1-\theta)Z+\theta Y)\|_2^2|\theta\right]\right]\\
&=\Ebb\left[\Ebb\left[\|Y-((1-\theta)Z+\theta Y)\|_2^2|\theta=1\right]\right]P(\theta=1)\\
&~~~+\Ebb\left[\Ebb\left[\|Y-((1-\theta)Z+\theta Y)\|_2^2|\theta=0\right]\right]P(\theta=0)\\
&=\Ebb[\|Y-Z\|_2^2|\theta=0]P(\theta=0)\\
&=\varepsilon \Ebb\|Y-Z\|_2^2\\
&\leq\mathcal{O}(p\varepsilon).
\end{align*}
Then we have
\begin{align}\label{W2mu}
W_2(\mu,\mu_{\varepsilon})\leq \mathcal{O}(\sqrt{p\varepsilon}).
\end{align}
Combining (\ref{ERRR1}) with (\ref{W2mu}), it yields  that
\begin{align}\label{error1}
&W_2(Law(\widetilde{Y}_{t_K}^{\varepsilon}),\mu)\notag\\
&~~\leq\mathcal{O}(\sqrt{p\varepsilon})+
\mathcal{O}(\sqrt{ps})+\mathcal{O}\left(\frac{\sqrt{p}(C_p)^2}{m^{1/4}\varepsilon^2}\right)+\mathcal{O}\left(\frac{(C_p)^2}{\sqrt{\log(m)}\varepsilon^2}\right)
+\mathcal{O}\left(\frac{\sqrt{p}C_p}{\sqrt{\log(m)}\varepsilon}\right).
\end{align}
Set $\varepsilon=(\log(m))^{-1/5}$ in (\ref{error1}),  then there exist one constant $\widetilde{C}_p$ depending on $p$ such that
\begin{align*}
W_2(Law(\widetilde{Y}_{t_K}^{\varepsilon}),\mu)
\leq\widetilde{C}_p\cdot\mathcal{O}\left(\frac{1}{(\log(m))^{1/10}}\right)+ \mathcal{O}(\sqrt{ps}).
\end{align*}

Moreover, if $f$ has the finite upper bound, then similar to the proof of Theorem \ref{th2} and by (\ref{W2mu}) and Lemma \ref{lemma61}, we have
\begin{align}\label{error2}
W_2(Law(\widetilde{Y}_{t_K}^{\varepsilon}),\mu)
\leq \mathcal{O}(\sqrt{p\varepsilon})+\mathcal{O}(\sqrt{ps})+
\mathcal{O}\left(\frac{\sqrt{p}(C_p)^2}{\sqrt{m}\varepsilon^2}\right)
+\mathcal{O}\left(\frac{\sqrt{p}C_p}{\sqrt{m}\varepsilon}\right).
\end{align}
Set $\varepsilon=m^{-1/5}$ in (\ref{error2}), then there exists one constant $\widetilde{C}_p$ depending on $p$ such that
\begin{align*}
W_2(Law(\widetilde{Y}_{t_K}^{\varepsilon}),\mu)
\leq\widetilde{C}_p\cdot\mathcal{O}\left(\frac{1}{m^{1/10}}\right)+ \mathcal{O}(\sqrt{ps}).
\end{align*}

\end{proof}

\end{appendix}

\bibliographystyle{siam}
\bibliography{ref_bib}

\end{document}